Plant scientists developed different varieties of corns that have a rich content of lysine which is a nutritious animal feed. A group of chicks were given this food to test the quality. The distribution of the weight gains (in grams) of these chicks are shown below:


Weight gains (in grams) Frequency

318 - 335 4

336 - 353 5

354 - 371 2

372 - 389 3

390 – 407 2

408 – 425 3

426 - 443 1


Find:

(a) the mean weight gains

(b) the median

(c) the variance for the above frequency intervals

(d) the standard deviation


1
Expert's answer
2021-03-31T09:25:55-0400

(a) Let's find the mean "\overline x = \frac{{\sum {x_i^*{n_i}} }}{n}" . Build a calculation table:

"\begin{matrix}
{interval}&{x_i^*,\,\,middle{\rm{ }}\,of{\rm{ }}\,interval}&{{n_i}}&{x_i^*{n_i}}\\
{318 - 335}&{326.5}&4&{1306}\\
{336 - 353}&{344.5}&5&{1722.5}\\
{354 - 371}&{362.5}&2&{725}\\
{372 - 389}&{380.5}&3&{1141.5}\\
{390 - 407}&{398.5}&2&{797}\\
{408 - 425}&{416.5}&3&{1249.5}\\
{426 - 431}&{434.5}&1&{434.5}\\
{sum}&{}&{20}&{7376}
\end{matrix}"

Then "\overline x = \frac{{7376}}{{20}} = {\rm{368}}{\rm{.8}}"

Answer: "\overline x = {\rm{368}}{\rm{.8}}"

(b) Let's make a table of accumulated frequencies "{f_i}" :

"\begin{matrix}
{interval}&{{n_i}}&{{f_i}}&{}\\
{318 - 335}&4&4&{}\\
{336 - 353}&5&9&{}\\
{354 - 371}&2&{11}&{}\\
{372 - 389}&3&{14}&{}\\
{390 - 407}&2&{16}&{}\\
{408 - 425}&3&{19}&{}\\
{426 - 431}&1&{20}&{}
\end{matrix}"

Then median interval is 354-371. Then median is

"Me = {x_{Me}} + i\frac{{\frac{{\sum f }}{2} - {f_{Me - 1}}}}{{{f_{Me}}}} = 354 + 17 \cdot \frac{{\frac{{20}}{2} - 9}}{{11}} \approx 355.55"

Answer: "Me \approx 355.55"

(c) Let's find the variance "D = \frac{{\sum {{{\left( {x_i^* - \overline x } \right)}^2}{n_i}} }}{n}" . Build a calculation table:

"\begin{matrix}
{interval}&{x_i^*,\,\,middle\,{\rm{ }}of{\rm{ }} \,interval}&{{n_i}}&{{{\left( {x_i^* - \overline x } \right)}^2}}&{{{\left( {x_i^* - \overline x } \right)}^2}{n_i}}\\
{318 - 335}&{326.5}&4&{1789.29}&{7157.16}\\
{336 - 353}&{344.5}&5&{590.49}&{2952.45}\\
{354 - 371}&{362.5}&2&{39.69}&{79.38}\\
{372 - 389}&{380.5}&3&{136.89}&{410.67}\\
{390 - 407}&{398.5}&2&{882.09}&{1764.18}\\
{408 - 425}&{416.5}&3&{2275.29}&{6825.87}\\
{426 - 431}&{434.5}&1&{4316.49}&{4316.49}\\
{sum}&{}&{20}&{}&{23506.2}
\end{matrix}"

Then "D = \frac{{23506.2}}{{20}} = {\rm{1175}}{\rm{.31}}"

Answer: "D = {\rm{1175}}{\rm{.31}}"

(d) Let's find the standard deviation: "\sigma {\rm{ = }}\sqrt D = \sqrt {{\rm{1175}}{\rm{.31}}} \approx 34.28" .

Answer: "\sigma {\rm{ }} \approx 34.28"